# Self-adjoint traces on the Pedersen ideal of $\mathrm{C}^{\ast}$-algebras

### v1

@@ 6. Affine functions on positive traces
The map is clearly continuous and maps the positive cone to the positive cone, and it is injective by Jordan decompositions (as in Proposition 6.1). It remains to show that the inverse map on the image is continuous.
@@
Notice that by Theorem 4.3 the embedding $T_{\mathbb{R}}(A)^{\ast}\to\mathrm{Aff}_{0}T_{+}(A)$ is surjective and thus an isomorphism if and only if the Trace Question (Question A) has a positive answer for $A$. The Trace Question is therefore essentially a question of whether the information of $T_{\mathbb{R}}(A)$ can be recovered from $T_{+}(A)$.
@@ 7. Traces on groupoid $\mathrm{C}^{\ast}$-algebras
An étale groupoid $\mathcal{G}$ is a topological groupoid such that the range map $r\colon\mathcal{G}\to\mathcal{G}^{0}$ (and thus also the source map) is a local homeomorphism, and we assume that the unit space $\mathcal{G}^{0}$ is locally compact and Hausdorff. By a twist over $\mathcal{G}$ we mean a Fell line bundle $\mathcal{L}\to\mathcal{G}$. We write $(\mathcal{G},\mathcal{L})$ for the resulting twisted étale groupoid. For an open bisection $U\subseteq\mathcal{G}$, we write $\Gamma_{c}(U,\mathcal{L})$ for the space of compactly supported continuous sections $U\to\mathcal{L}$. We include $\Gamma_{c}(U,\mathcal{L})$ into the space of bounded sections $\mathcal{G}\to\mathcal{L}$ by setting the value to be 00 outside $U$, and set $\Gamma_{c}(\mathcal{G},\mathcal{L})$ to be the span of all these sections over open bisections $U$. When $\mathcal{G}$ is Hausdorff, this is the space of compactly supported sections, but for non-Hausdorff groupoids we warn the reader that these sections are not necessarily continuous. We identify the space of sections $\Gamma_{c}(\mathcal{G}^{0},\mathcal{L})$ supported on the unit space with $C_{c}(\mathcal{G}^{0})$ as the restriction of $\mathcal{L}$ to $\mathcal{G}^{0}$ is the trivial bundle whose fibres have the structure of the $\mathrm{C}^{\ast}$-algebra $\mathbb{C}$. We refer to [EP22] for an introduction to twisted étale groupoids from the line bundle perspective and their $\mathrm{C}^{\ast}$-algebras.
Convolution defines a *-algebra structure on $\Gamma_{c}(\mathcal{G},\mathcal{L})$. Given a faithful *-representation $\pi\colon\Gamma_{c}(\mathcal{G},\mathcal{L})\to\mathcal{B}(\mathcal{H})$ we write $\|\cdot\|_{\pi}$ for the induced $\mathrm{C}^{\ast}$-norm on $\Gamma_{c}(\mathcal{G},\mathcal{L})$ and we denote the completion by $\mathrm{C}^{\ast}_{\pi}(\mathcal{G},\mathcal{L})$. For the trivial Fell line bundle =×$\mathcal{L}=\mathbb{C}\times\mathcal{G}\to\mathcal{G}$ we recover the usual untwisted *-algebras and write $C_{c}(\mathcal{G})$ and $\mathrm{C}^{\ast}_{\pi}(\mathcal{G})$. The reduced $\mathrm{C}^{\ast}$-algebra $\mathrm{C}^{\ast}*_{\lambda}(\mathcal{G},\mathcal{L})$ is the $\mathrm{C}^{\ast}$-algebra defined by the left regular representation $\lambda\colon\Gamma_{c}(\mathcal{G},\mathcal{L})\to\mathcal{B}(\bigoplus_{x\in G^{0}}\ell^{2}(\mathcal{G}_{x},\mathcal{L}))$.
@@
If $(\mathcal{G},\mathcal{L})$ is a twisted étale groupoid and $Y\subseteq\mathcal{G}^{0}$ is open, we let $\Gamma_{c}(\mathcal{G},\mathcal{L})|_{Y}$ denote the $\ast$-subalgebra of all $f\in\Gamma_{c}(\mathcal{G},\mathcal{L})$ for which $f=0$ outside of $\mathcal{G}^{Y}_{Y}$. The following describes $T_{\mathbb{R}}(\mathrm{C}^{\ast}_{\pi}(\mathcal{G},\mathcal{L}))$ as an ordered vector space in terms of traces on $\Gamma_{c}(\mathcal{G},\mathcal{L})$. We note that one could also describe the topology using Proposition 5.3, but that the topology is not obviously natural in this case.
@@
For Hausdorff groupoids we have a canonical conditional expectation $E\colon\mathrm{C}^{\ast}_{\pi}(\mathcal{G},\mathcal{L})\to C_{0}(\mathcal{G})$ extending the map $\Gamma_{c}(\mathcal{G},\mathcal{L})\to C_{c}(\mathcal{G}^{0})$ given by $f\mapsto f|_{\mathcal{G}^{0}}$. This is often used to induce bounded traces on $\mathrm{C}^{\ast}_{\pi}(\mathcal{G},\mathcal{L})$ coming from bounded $\mathcal{G}$-invariant measures on $\mathcal{G}^{0}$. However, $E$ does not map $\mathrm{Ped}(\mathrm{C}^{\ast}_{\pi}(\mathcal{G},\mathcal{L}))$ into $C_{c}(\mathcal{G}^{0})=\mathrm{Ped}(C_{0}(\mathcal{G}^{0}))$ in general, so this approach cannot a priori be generalised to unbounded traces.
For instance, if $(C_{0}(\mathcal{G}^{0}),\mathrm{C}^{\ast}_{\pi}(\mathcal{G},\mathcal{L}))=(C_{0}(\mathbb{N}),\mathcal{K}(\ell^{2}(\mathbb{N})))$ with standard matrix units $e_{j,k}\in\mathcal{K}(\ell^{2}(\mathbb{N}))$, then the rank one projection $p=\sum_{j,k=1}^{\infty}\tfrac{1}{jk}e_{j,k}$ is in $\mathrm{Ped}(\mathcal{K}(\ell^{2}(\mathbb{N})))$ (since it is finite rank), but $E(p)=\sum_{j=1}^{\infty}j^{-2}e_{j,j}\in C_{0}(\mathbb{N})$ which does not have compact support.
@@
As $C_{c}(\mathcal{G}^{0})$ generates $\mathrm{C}^{\ast}_{\pi}(\mathcal{G},\mathcal{L})$ as a two-sided closed ideal, we may by Lemma 5.2 find $m\in\mathbb{N}$, $g_{1},\dots,g_{m}\in C_{c}(\mathcal{G}^{0})$ and $c_{1},\dots,c_{m}\in\mathrm{C}^{\ast}_{\pi}(\mathcal{G},\mathcal{L})$ such that $a=\sum_{j=1}^{m}c_{j}^{\ast}g_{j}^{\ast}g_{j}c_{j}$. By considering each term in this sum individually, we may assume without loss of generality that $a=cg^{\ast}gc$ for some $g\in C_{c}(\mathcal{G}^{0})$ and $c\in\mathrm{C}^{\ast}_{\pi}(\mathcal{G},\mathcal{L})$.
@@
The final part follows from Corollary 7.5 since principal groupoids are automatically Hausdorff. ∎

### v2

@@ 6. Affine functions on positive traces
The map is clearly continuous and maps the positive cone to the positive cone, and it is injective by Jordan decompositions (as in the proof of Proposition 6.1). It remains to show that the inverse map on the image is continuous.
@@
Notice that by Theorem 4.3, surjectivity of the embedding $T_{\mathbb{R}}(A)^{\ast}\to\mathrm{Aff}_{0}T_{+}(A)$ (and thus that it is an isomorphism) is equivalent to a positive answer to the Trace Question (Question A) for $A$. The Trace Question is therefore essentially a question of whether the information of $T_{\mathbb{R}}(A)$ can be recovered from $T_{+}(A)$.
@@ 7. Traces on groupoid $\mathrm{C}^{\ast}$-algebras
An étale groupoid $\mathcal{G}$ is a topological groupoid such that the range map $r\colon\mathcal{G}\to\mathcal{G}^{0}$ (and thus also the source map) is a local homeomorphism, and we assume that the unit space $\mathcal{G}^{0}$ is locally compact and Hausdorff. By a twist over $\mathcal{G}$ we mean a Fell line bundle $\mathcal{L}\to\mathcal{G}$, and we call the pair $(\mathcal{G},\mathcal{L})$ a twisted étale groupoid. For an open bisection $U\subseteq\mathcal{G}$, we write $\Gamma_{c}(U,\mathcal{L})$ for the space of compactly supported continuous sections $U\to\mathcal{L}$. We include $\Gamma_{c}(U,\mathcal{L})$ into the space of bounded sections $\mathcal{G}\to\mathcal{L}$ by setting the value to be 00 outside $U$, and set $\Gamma_{c}(\mathcal{G},\mathcal{L})$ to be the span of all these sections over open bisections $U$. When $\mathcal{G}$ is Hausdorff, this is the space of compactly supported sections, but for non-Hausdorff groupoids we warn the reader that these sections are not necessarily continuous. We identify the space of sections $\Gamma_{c}(\mathcal{G}^{0},\mathcal{L})$ supported on the unit space with $C_{c}(\mathcal{G}^{0})$ as the restriction of $\mathcal{L}$ to $\mathcal{G}^{0}$ is the trivial bundle whose fibres have the structure of the $\mathrm{C}^{\ast}$-algebra $\mathbb{C}$. We refer to [EP22] for an introduction to twisted étale groupoids from the line bundle perspective and their $\mathrm{C}^{\ast}$-algebras.
Convolution defines a *-algebra structure on $\Gamma_{c}(\mathcal{G},\mathcal{L})$. Given a faithful *-representation $\pi\colon\Gamma_{c}(\mathcal{G},\mathcal{L})\to\mathcal{B}(\mathcal{H})$ we write $\|\cdot\|_{\pi}$ for the induced $\mathrm{C}^{\ast}$-norm on $\Gamma_{c}(\mathcal{G},\mathcal{L})$ and we denote the completion by $\mathrm{C}^{\ast}_{\pi}(\mathcal{G},\mathcal{L})$. For the trivial Fell line bundle =×$\mathcal{L}=\mathbb{C}\times\mathcal{G}\to\mathcal{G}$ we recover the usual untwisted *-algebras and write $C_{c}(\mathcal{G})$ and $\mathrm{C}^{\ast}_{\pi}(\mathcal{G})$. The reduced $\mathrm{C}^{\ast}$-algebra $\mathrm{C}^{\ast}_{\lambda}(\mathcal{G},\mathcal{L})$ is the $\mathrm{C}^{\ast}$-algebra defined by the left regular representation $\lambda\colon\Gamma_{c}(\mathcal{G},\mathcal{L})\to\mathcal{B}(\bigoplus_{x\in G^{0}}\ell^{2}(\mathcal{G}_{x},\mathcal{L}))$.
@@
If $(\mathcal{G},\mathcal{L})$ is a twisted étale groupoid and $Y\subseteq\mathcal{G}^{0}$ is open, we let $\Gamma_{c}(\mathcal{G},\mathcal{L})|_{Y}$ denote the $\ast$-subalgebra of all $f\in\Gamma_{c}(\mathcal{G},\mathcal{L})$ for which $f=0$ outside of $\mathcal{G}^{Y}_{Y}:=\{g\in\mathcal{G}:r(g),s(g)\in Y\}$. The following describes $T_{\mathbb{R}}(\mathrm{C}^{\ast}_{\pi}(\mathcal{G},\mathcal{L}))$ as an ordered vector space in terms of traces on $\Gamma_{c}(\mathcal{G},\mathcal{L})$. We note that one could also describe the topology using Proposition 5.3, but that the topology is not obviously natural in this case.
@@
For Hausdorff groupoids we have a canonical conditional expectation $E\colon\mathrm{C}^{\ast}_{\pi}(\mathcal{G},\mathcal{L})\to C_{0}(\mathcal{G}^{0})$ extending the map $\Gamma_{c}(\mathcal{G},\mathcal{L})\to C_{c}(\mathcal{G}^{0})$ given by $f\mapsto f|_{\mathcal{G}^{0}}$. This is often used to induce bounded traces on $\mathrm{C}^{\ast}_{\pi}(\mathcal{G},\mathcal{L})$ coming from bounded $\mathcal{G}$-invariant measures on $\mathcal{G}^{0}$. However, $E$ does not map $\mathrm{Ped}(\mathrm{C}^{\ast}_{\pi}(\mathcal{G},\mathcal{L}))$ into $C_{c}(\mathcal{G}^{0})=\mathrm{Ped}(C_{0}(\mathcal{G}^{0}))$ in general, so this approach cannot a priori be generalised to unbounded traces.
For instance, if $(C_{0}(\mathcal{G}^{0}),\mathrm{C}^{\ast}_{\pi}(\mathcal{G},\mathcal{L}))=(C_{0}(\mathbb{N}),\mathcal{K}(\ell^{2}(\mathbb{N})))$ with standard matrix units $e_{j,k}\in\mathcal{K}(\ell^{2}(\mathbb{N}))$, then the rank one projection $p=\sum_{j,k=1}^{\infty}\tfrac{\sqrt{6}}{jk\pi}e_{j,k}$ is in $\mathrm{Ped}(\mathcal{K}(\ell^{2}(\mathbb{N})))$ (since it is finite rank), but $E(p)=\sum_{j=1}^{\infty}\tfrac{\sqrt{6}}{j^{2}\pi}e_{j,j}\in C_{0}(\mathbb{N})$ which does not have compact support.
@@
As $C_{c}(\mathcal{G}^{0})$ generates $\mathrm{C}^{\ast}_{\pi}(\mathcal{G},\mathcal{L})$ as a two-sided closed ideal, we may by Lemma 5.2 find $m\in\mathbb{N}$, $g_{1},\dots,g_{m}\in C_{c}(\mathcal{G}^{0})$ and $c_{1},\dots,c_{m}\in\mathrm{C}^{\ast}_{\pi}(\mathcal{G},\mathcal{L})$ such that $a=\sum_{j=1}^{m}c_{j}^{\ast}g_{j}^{\ast}g_{j}c_{j}$. By considering each term in this sum individually, we may assume without loss of generality that $a=c^{\ast}g^{\ast}gc$ for some $g\in C_{c}(\mathcal{G}^{0})$ and $c\in\mathrm{C}^{\ast}_{\pi}(\mathcal{G},\mathcal{L})$.
@@
The final part follows from Corollary 7.5 since principal groupoids are automatically Hausdorff. [^11] If $g_{1},g_{2}\in\mathcal{G}$ are distinct, principality implies that either $s(g_{1})\neq s(g_{2})$ or $r(g_{1})\neq r(g_{2})$, after which Hausdorffness of $\mathcal{G}^{0}$ can be used to separate $g_{1}$ and $g_{2}$ by open sets in $\mathcal{G}$.